# Machine Unlearning for Image-to-Image Generative Models

### v1

@@ 4.1 Experimental Setup
Table 1: Results of cropping $\times 8$ 8 patches at the center of the image, where each patch is $\times 16$ 16 pixels. $\uparrow$’ means higher is better and $\downarrow$’ means lower is better. $R$ and $F$ account for the retain set and forget set, respectively.‘Proxy $\mathcal{D}_{R}$’ means that we use the images from other classes as a substitute of the real retain set to do the unlearning (cf. Section 4.3).
@@
Table 2: Results of cropping $\times 4$ 4 patches at the center of the image, where each patch is $\times 16$ 16 pixels. $\uparrow$’ means higher is better and $\downarrow$’ means lower is better. $R$ and $F$ account for the retain set and forget set, respectively. “Proxy $\mathcal{D}_{R}$” means that we use the images from other classes as a substitute of the real retain set to do the unlearning (cf. Section 4.3).
@@
Figure 3: Results of cropping $\times 8$ 8 patches at the center of the image on diffusion models, where each patch is $\times 16$ 16 pixels. Our method has negligible-to-slight performance degradation on diverse I2I generative models and multiple generative tasks. (cf. Appendix D and E.1).
@@
Table 3: Ablation study of $\alpha$’s values. We test the performance of cropping $\times 8$ 8 patches at the center of the image. As shown, $\alpha=0.25$ achieves a good balance between the preserving the performance on retain set, while remove the information on forget sets across these two models.
@@ 4.4 Ablation Study
For the ablation study, we test the results of cropping patches at the center of the image under various setups, where each patch is $\times 16$ 16 pixels.
@@ Appendix A Infinitely many probability with infinite KL-divergence
[…]
@@
[…]
@@ C.3 Training Hyper-parameters
Patch Size For all of the models we evaluate in this paper (including diffusion models, VQ-GAN and MAE), the networks are vision transformer (Dosovitskiy et al., 2021) based architecture. We set the size of each patch as $\times 16$ 16 pixels for all experiments. For example, cropping $\times 8$ 8 patches means removing $\times 128$ 128 pixels. We set $\alpha=0.25$ (cf. Eq. (10)).
@@ D.1 Robustness to Retain Samples Availability
[…]
@@
Figure D.7: Diffusion models: cropping $\times 4$ 4 patches at the center of the image, where each patch is $\times 16$ 16 pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise. Overall, we outperform all other baseline methods.
@@
Figure D.8: Diffusion models: cropping $\times 8$ 8 patches at the center of the image, where each patch is $\times 16$ 16 pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise. Overall, we outperform all other baseline methods.
@@ D.2 Diffusion
We provide the visualization results under center crop size of $\times 4$ 4 patches in Fig. D.7. We also provide more visualization results under center crop size of $\times 8$ 8 patches in Fig. D.8. As shown in Fig. D.7 and Fig. D.8, out approach can generate very high quality images on retain set while filling some random noise on the forget set. As a contrast, the other baselines methods struggle on the performance drop on the retain set. Random Encoder cannot even forget the information on the forget set.
@@
Figure D.9: VQ-GAN: cropping $\times 4$ 4 patches at the center of the image, where each patch is $\times 16$ 16 pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise.
@@
Figure D.10: VQ-GAN: cropping $\times 8$ 8 patches at the center of the image, where each patch is $\times 16$ 16 pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise.
@@ D.3 VQ-GAN
We provide the visualization results under center crop size of $\times 4$ 4 patches in Fig. D.9. We also provide more visualization results under center crop size of $\times 8$ 8 patches in Fig. D.10. As shown in Fig. D.9 and Fig. D.10, out approach can generate very high quality images on retain set while fill some random noise on the forget set. Retain Label cannot even forget the information on the forget set. Moreover, as shown in Fig. D.9, there are two type of fishes; one is in forget set and another one is in retain set (the fourth and fifth rows). Our approach can handle this subtle difference among difference classes.
@@ D.4 MAE
Figure D.11: The quality of reconstructed images given random masked images by MAE under varying random masking ratio; e.g., masking 128128128 out of 256 patches means 50% cropping ratio. We compare the original model and the unlearned model by our approach. As shown, the performance of these two models on the retain set are almost identical, while the unlearned model has a significant performance drop on the forget set. See Fig. D.4 for some examples of generated images.
@@ E.1 Varying Cropping Pattern and Ratio
Figure E.20: Ablation study of noise type. We test the performance under varying ratios of central cropping; e.g., cropping $\times 8$ 8 out of 256 patches means a 25% cropping ratio. For VQ-GAN and MAE, Gaussian noise achieves better forgetting in the forget set in general (i.e., it achieves higher FID, lower IS and lower CLIP on forget set).
@@ E.2 Cross Validation of Generated Images
  • Second, given the center-cropped input images (cropping central $\times 8$ 8 patches), we then use this unlearned model to generate the images on both forget set and retain set (i.e., image uncropping/inpainting). Here we call them First generated images.
  • Given these First generated images, we only keep the reconstructed central $\times 8$ 8 patches (cropping the outer ones) as the new input to the original VQ-GAN (i.e., the model before unlearning) and get the newly generated images (here we call them Second generated images). We then evaluate the quality of Second generated images for both forget set and retain set and report the results.

### v2

@@ 4.1 Experimental Setup
Table 1: Results of cropping $8\times 8$ patches at the center of the image, where each patch is $16\times 16$ pixels. $\uparrow$’ means higher is better and $\downarrow$’ means lower is better. $R$ and $F$ account for the retain set and forget set, respectively.‘Proxy $\mathcal{D}_{R}$’ means that we use the images from other classes as a substitute of the real retain set to do the unlearning (cf. Section 4.3).
@@
Table 2: Results of cropping $4\times 4$ patches at the center of the image, where each patch is $16\times 16$ pixels. $\uparrow$’ means higher is better and $\downarrow$’ means lower is better. $R$ and $F$ account for the retain set and forget set, respectively. “Proxy $\mathcal{D}_{R}$” means that we use the images from other classes as a substitute of the real retain set to do the unlearning (cf. Section 4.3).
@@
Figure 3: Results of cropping $8\times 8$ patches at the center of the image on diffusion models, where each patch is $16\times 16$ pixels. Our method has negligible-to-slight performance degradation on diverse I2I generative models and multiple generative tasks. (cf. Appendix D and E.1).
@@
Table 3: Ablation study of $\alpha$’s values. We test the performance of cropping $8\times 8$ patches at the center of the image. As shown, $\alpha=0.25$ achieves a good balance between the preserving the performance on retain set, while remove the information on forget sets across these two models.
@@ 4.4 Ablation Study
For the ablation study, we test the results of cropping patches at the center of the image under various setups, where each patch is $16\times 16$ pixels.
@@ Appendix A Infinitely many probability with infinite KL-divergence
where $j$ satisfies $P(X=j)>0$ and $k$ sastifies $j\neq k$. Clearly, $10<=Q(X=i)<$ 1 and $\sum_{i=1}^{N}Q(X=i)=1$. Therefore, $Q(X)$ is a valid probability distribution.
@@
We note that $\kappa$ can be any positive real number; hence, we can obtain infinitely many $Q(X)$ by varying the value of $\kappa$. Hence, there are infinitely many $Q$ that have a positively infinite value of KL-divergence with $P$. In general, one can replace the set $[N]$ with any discrete set. This completes our proof. $\Box$
@@ C.3 Training Hyper-parameters
Patch Size For all of the models we evaluate in this paper (including diffusion models, VQ-GAN and MAE), the networks are vision transformer (Dosovitskiy et al., 2021) based architecture. We set the size of each patch as $16\times 16$ pixels for all experiments. For example, cropping $8\times 8$ patches means removing $128\times 128$ pixels. We set $\alpha=0.25$ (cf. Eq. (10)).
@@ D.1 Robustness to Retain Samples Availability
We next vary the number of images per class for the retain set within the range of $\{1,5,10,20,25,50\}$ and compare the performance of the unlearned models under different values. To balance the number of forget sample and retain samples, we over sample the retain samples. For example, if we have 5 image per class for the retain set, we sample these images 20 times (since $100/5=20$).
@@
Figure D.7: Diffusion models: cropping $4\times 4$ patches at the center of the image, where each patch is $16\times 16$ pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise. Overall, we outperform all other baseline methods.
@@
Figure D.8: Diffusion models: cropping $8\times 8$ patches at the center of the image, where each patch is $16\times 16$ pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise. Overall, we outperform all other baseline methods.
@@ D.2 Diffusion
We provide the visualization results under center crop size of $4\times 4$ patches in Fig. D.7. We also provide more visualization results under center crop size of $8\times 8$ patches in Fig. D.8. As shown in Fig. D.7 and Fig. D.8, out approach can generate very high quality images on retain set while filling some random noise on the forget set. As a contrast, the other baselines methods struggle on the performance drop on the retain set. Random Encoder cannot even forget the information on the forget set.
@@
Figure D.9: VQ-GAN: cropping $4\times 4$ patches at the center of the image, where each patch is $16\times 16$ pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise.
@@
Figure D.10: VQ-GAN: cropping $8\times 8$ patches at the center of the image, where each patch is $16\times 16$ pixels. As shown, our approach has almost identical performance as the original model (i.e., before unlearning); the generated images on forget set are some random noise.
@@ D.3 VQ-GAN
We provide the visualization results under center crop size of $4\times 4$ patches in Fig. D.9. We also provide more visualization results under center crop size of $8\times 8$ patches in Fig. D.10. As shown in Fig. D.9 and Fig. D.10, out approach can generate very high quality images on retain set while fill some random noise on the forget set. Retain Label cannot even forget the information on the forget set. Moreover, as shown in Fig. D.9, there are two type of fishes; one is in forget set and another one is in retain set (the fourth and fifth rows). Our approach can handle this subtle difference among difference classes.
@@ D.4 MAE
Figure D.11: The quality of reconstructed images given random masked images by MAE under varying random masking ratio; e.g., masking $128$ out of 256 patches means 50% cropping ratio. We compare the original model and the unlearned model by our approach. As shown, the performance of these two models on the retain set are almost identical, while the unlearned model has a significant performance drop on the forget set. See Fig. D.4 for some examples of generated images.
@@ E.1 Varying Cropping Pattern and Ratio
Figure E.20: Ablation study of noise type. We test the performance under varying ratios of central cropping; e.g., cropping $8\times 8$ out of 256 patches means a 25% cropping ratio. For VQ-GAN and MAE, Gaussian noise achieves better forgetting in the forget set in general (i.e., it achieves higher FID, lower IS and lower CLIP on forget set).
@@ E.2 Cross Validation of Generated Images
  • Second, given the center-cropped input images (cropping central $8\times 8$ patches), we then use this unlearned model to generate the images on both forget set and retain set (i.e., image uncropping/inpainting). Here we call them First generated images.
  • Given these First generated images, we only keep the reconstructed central $8\times 8$ patches (cropping the outer ones) as the new input to the original VQ-GAN (i.e., the model before unlearning) and get the newly generated images (here we call them Second generated images). We then evaluate the quality of Second generated images for both forget set and retain set and report the results.